OpenStudy (anonymous):

The function f(x) passes through the point (-3, 3). If g(x)=f(x-3)+4, then which one of these points, if any, must lie on the graph of g(x)?

OpenStudy (anonymous):

\[x-3=-3\\ x=0\]

OpenStudy (anonymous):

so if \[f(-3)=3\] then \[g(0)=f(0-3)+4=3+4\]

OpenStudy (anonymous):

wow thats a lot to process ok so then what

OpenStudy (anonymous):

add \(3+4\)

OpenStudy (anonymous):

it does look confusing, but one thing that should not be confusing is that if \[(-3,3)\] is on the graph of \(f\) then that measn \[f(-3)=3\] right?

OpenStudy (anonymous):

okay so that means it equals 7.. what then

OpenStudy (anonymous):

lets back up

OpenStudy (anonymous):

\[g(x)=f(x-3)+4\]

OpenStudy (anonymous):

you only know what \(f(-3)\) is

OpenStudy (anonymous):

so the only way to know a value for \(g\) is if the input is \(-3\) which will be the case if \(x=0\) since \[0-3=-3\]

OpenStudy (anonymous):

if \(x=0\) then \(g(0)=f(0-3)+4=f(-3)+4=7\)

OpenStudy (anonymous):

that means, on the graph of \(g\) is the point \((0,7)\)

OpenStudy (anonymous):

okay so how would you do this next one.. i kinda understand. the function f(x) passes through (4,-2). If g(x)=-f(x), which one of these points must lie on graph of g(x)

OpenStudy (anonymous):

so if \(f(4)=-2\) the \(g(4)=-f(4)=-(-2)=2\)

OpenStudy (anonymous):

which is a long winded way of saying if \((4,-2)\) is on the graph of \(f\) then \((4,2)\) is on the graph of \(-f\)

OpenStudy (anonymous):

wow you are a god. okay ill probably have another question soon dont leave meh

OpenStudy (anonymous):

ok dear i will stay

OpenStudy (anonymous):

not a god however, just a satellite